43) Three resistors, 25, 45, and 75.2, are connected in a series, and a 0.51-A current passes through them. What are (a) the equivalent resistance and (b) the potential difference across the three resistors?
Answer #1

a)

R = 25 + 45 + 75 = 145 ohm

b)

V = i R = 0.51* 145

V = 73.95 V

========

Comment in case any doubt, will reply for sure. Goodluck